Business Solutions's second-quarter 2018 fixed budget performance report for its computer furniture operations follows. The $168,990 budgeted expenses include $111,200 in variable expenses for desks and $20,790 in variable expenses for chairs, as well as $37,000 fixed expenses. The actual expenses include $38,000 fixed expenses. List fixed and variable expenses separately.

Fixed Budget Actual Results Variances
Desk sales (in units) 139 145
Chair sales (in units) 63 71
Desk sales $ 179,310 $ 185,600 $ 6,290 F
Chair sales 34,020 39,405 5,385 F
Total expenses 168,990 177,810 8,820 U
Income from operations $ 44,340 $ 47,195 $ 2,855 F


Prepare a flexible budget performance report that shows any variances between budgeted results and actual results.
Answer #1
Flexible Budget Actual Results Variance Fav./Unf.
Desk sales 187050 185600 1450 Unfavorable
Chair sales 38340 39405 1065 Favorable
Variable expenses 139430 139810 380 Unfavorable
Contribution margin 85960 85195 765 Unfavorable
Fixed expenses 37000 38000 1000 Unfavorable
Income from operations 48960 47195 1765 Unfavorable
Workings:
Flexible Budget
Desk sales 187050 =179310/139*145
Chair sales 38340 =34020/63*71
Variable expenses 139430 =(111200/139*145)+(20790/63*71)